# Noise-Aware Differentially Private Regression via Meta-Learning

### v1

@@ 1 Introduction
Neural Processes. Our method is based on neural processes (NPs; Garnelo et al., 2018a), a model which leverages the flexibility of neural networks to produce well-calibrated predictions in the meta-learning setting. The parameters of the NP can be meta-trained to generalise to unseen datasets, while being much faster than gradient-based fine-tuning alternatives (Finn et al., 2017).
@@
  • We improve upon the functional mechanism of Hall et al. (2013) by leveraging Gaussian DP theory (Dong et al., 2022), showing that context set privacy can be protected with smaller amounts of noise (at least 25% lower standard deviation in the settings considered in Figure 4). We incorporate these improvements into DPConvCNP, but note that they are also of independent interest.
@@ 2 Related Work
Training deep learning models on public proxy datasets and then fine-tuning with DP on private data is becoming increasingly common in computer vision and natural language processing applications (Yu et al., 2021; Li et al., 2022; De et al., 2022; Tobaben et al., 2023). However, these approaches rely on the availability of very large non-sensitive datasets. Because these datasets would likely need to be scraped from the internet, it is unclear whether they are actually non-sensitive (Tramèr et al., 2022). On the other hand, other approaches study meta-learning with DP during meta-training (Li et al., 2020; Zhou and Bassily, 2022), but do not enforce privacy guarantees at meta-test time.
@@ 3.1 Meta-learning and Neural Processes
In CNPs, a standard choice which we also use here, is to let $\pi_{\phi,\theta}(\mathbf{x}^{(t)},D^{(c)})$ return a mean $\mu_{\phi,\theta}(x^{(t)},D^{(c)})$ and a variance $\sigma_{\phi,\theta}^{2}(x^{(t)},D^{(c)}),$ to parameterise a predictive that factorises across the target points $y^{(t)}|x^{(t)}\sim\mathcal{N}(\mu_{\phi,\theta}(x^{(t)},D^{(c)}),\sigma_{\phi,\theta}^{2}(x^{(t)},D^{(c)})).$ We note that our framework straightforwardly extends to more complicated $\pi_{\phi,\theta}(\mathbf{x}^{(t)},D^{(c)})$. To train a CNP to make accurate predictions, we can optimise a log-likelihood objective Garnelo et al. (2018a) such as
@@
  Input: Datasets $(D_{m})^{M}_{m=1}$, encoder $\texttt{enc}_{\phi}$, decoder $\texttt{dec}_{\theta},$ iterations $T$, optimiser opt
@@
  Input: Context $D^{(c)},\texttt{enc}_{\phi},\texttt{dec}_{\theta}.$
@@ 3.2 Differential Privacy
Neighbourhood definition domain-specific. We consider $D\in\mathbb{R}^{N\times d}$ with $N$ users and $d$ dimensions and the substitution neighbourhood relation $\sim_{S}$ where $D\sim_{S}D^{\prime}$ if $D$ and $D^{\prime}$ differ by one row.
@@
Properties of (G)DP. Differential privacy has several useful properties. First, post-processing immunity guarantees that post-processing the result of a DP algorithm does not cause privacy loss: {theorem}[Dwork and Roth 2014] Let $\mathcal{M}$ be an $(\epsilon,\delta)$-DP (or $\mu$-GDP) algorithm and let $f$ be any, possibly randomised, function. Then $f\circ\mathcal{M}$ is $(\epsilon,\delta)$-DP (or $\mu$-GDP). Composition of DP mechanisms refers to running multiple mechanisms on the same data. When each mechanism can depend on the outputs of the previous mechanisms, the composition is called adaptive. GDP is particularly appealing because it has a simple and tight composition formula: {theorem}[Dong et al. 2022] The adaptive composition of $T$ mechanisms that are $\mu_{i}$-GDP ($i=1,\dots,T$), is $\mu$-GDP with $\mu=\sqrt{\mu_{1}^{2}+\dotsb+\mu_{T}^{2}}.$ Gaussian mechanism. One of the central mechanisms to guarantee DP, is the Gaussian mechanism. This releases the output of a function $f$ with added Gaussian noise $\mathcal{M}(D)=f(D)+\mathcal{N}(0,\sigma^{2}I)$, where the variance $\sigma^{2}$ depends on the $l_{2}$-sensitivity of $f,$ defined as $\Delta=\sup_{D\sim D^{\prime}}||f(D)-f(D^{\prime})||_{2}.$theorem}[Dong et al. 2022] The Gaussian mechanism with variance $\sigma^{2}=\nicefrac{{\Delta^{2}}}{{\mu^{2}}}$ is $\mu$-GDP.
@@ 3.3 The Functional Mechanism
The functional mechanism with multiplier $c$ and sensitivity $\Delta$ is defined as $\mathcal{M}(D)=f_{D}+cg.$
@@ 4 Differential privacy for the ConvCNP
Now we turn to the main contributions of our work. First, we tighten the functional mechanism in Section 4.1 and then we build the functional mechanism into the ConvCNP in Section 4.2.
@@ 4.1 Improving the Functional Mechanism
Figure 4: Noise magnitude comparison for the classical functional mechanism of Hall et al. (2013), the RDP-based mechanism of Jiang et al. (2023) and our improved GDP-based mechanism. The line for Hall et al. cuts off at $\epsilon=1$ since their bound has only been proven for $\epsilon\leq 1$. We set $\Delta^{2}=10$ and $\delta=10^{-3}$, which are representative values from our experiments.
@@ 4.2 Differentially Private Convolutional CNP
which are the two quantities we want to release under DP. To achieve this, we must first determine the sensitivity of $r^{(d)}$ and $r^{(s)},$ as defined in Eq. 6. Recall that we use the substitution neighbourhood relation $\sim_{S}$ defined as $D^{(c)}_{1}\sim_{S}D^{(c)}_{2}$ if $D^{(c)}_{1}$ and $D^{(c)}_{2},$ differ in at most one row, i.e. by a single context point. Since the RBF $\psi$ is bounded above by $1,$ it can be shown (see Section A.4) that the squared $l_{2}$-sensitivity of $r^{(d)}$ is bounded above by $2,$ and this bound is tight. Unfortunately however, since the signal channel $r^{(s)}$ depends linearly on each $y^{(c)}_{n}$ (see Eq. 7), its sensitivity is unbounded. To address this, we clip each $y^{(c)}_{n}$ by a threshold $C,$ which is a standard way to ensure the sensitivity is bounded. With this modification we obtain the following tight sensitivities for $r^{(d)}$ and $r^{(s)}$: $\Delta_{\mathcal{H}}^{2}r^{(d)}=2$ and $\Delta_{\mathcal{H}}^{2}r^{(s)}=4C^{2}$. With these in place, we can state our privacy guarantee which forms the basis of the DPConvCNP. Post-processing immunity (Section 3.2) ensures that post-processing $r^{(s)}$ and $r^{(d)}$ with the ConvCNP decoder does not result in further privacy loss. {theorem} Let $g_{d}$ and $g_{s}$ be sample paths of two independent Gaussian processes having zero mean and covariance function $k,$ such that $0\leq k\leq C_{k}$ for some $C_{k}>0.$ Let $\Delta_{d}^{2}=2C_{k}$ and $\Delta_{s}^{2}=4C^{2}C_{k}.$ Then releasing $r^{(d)}+\sigma_{d}g_{d}$ and $r^{(s)}+\sigma_{s}g_{s}$ is $\mu$-GDP with $\mu=\sqrt{\nicefrac{{\Delta_{s}^{2}}}{{\sigma_{s}^{2}}}+\nicefrac{{\Delta_{d}^{2}}}{{\sigma_{d}^{2}}}}.$
@@ 4.3 Training the DPConvCNP
Privacy hyperparameters. By Section 3.2 and Section 4.2, each $(\epsilon,\delta)$-budget implies a $\mu$-budget, placing a constraint on the sensitivities and noise magnitudes, namely $\mu^{2}=\nicefrac{{\Delta_{s}^{2}}}{{\sigma_{s}^{2}}}+\nicefrac{{\Delta_{d}^{2}}}{{\sigma_{d}^{2}}}.$ Since $\psi$ is an RBF, $\Delta_{d}^{2}=2$ and $\Delta_{s}^{2}=4C^{2},$ and we need to specify $C,\sigma_{s}$ and $\sigma_{d}$, subject to this constraint. We introduce a variable $10<t<$ 1 and rewrite the constraint as $\sigma_{s}^{2}=\nicefrac{{4C^{2}}}{{t\mu^{2}}}$ and $\sigma_{d}^{2}=\nicefrac{{2}}{{(1-t)\mu^{2}}}$ allowing us to freely set $t$ and $C.$ One straightforward approach is to fix $t$ and $C$ to hand-picked values, but this is sub-optimal since the optimal values depend on $\mu,N,$ and the statistics of the data. Instead, we can make them adaptive, letting $t:\mathbb{R}^{+}\times\mathbb{N}\to(0,1)$ and $C:\mathbb{R}^{+}\times\mathbb{N}\to\mathbb{R}^{+}$ be learnable functions, e.g. neural networks $t(\mu,N)=\texttt{sig}(\text{NN}_{t}(\mu,N))$ and $C(\mu,N)=\text{exp}(\text{NN}_{C}(\mu,N))$ where sig is the sigmoid. These neural networks are meta-trained along with all other parameters of the DPConvCNP.
@@ 5 Experiments & Discussion
We conduct experiments on synthetic and a sim-to-real task with real data. We provide the exact experimental details in Appendix E.
DP-SGD baseline. Since, we are interested in the small-data regime, i.e. a few hundred of datapoints per task, we turn to Gaussian processes (GP; Rasmussen and Williams, 2006), the gold-standard model for well-calibrated predictions in this setting. To enforce DP, we make the GP variational (Titsias, 2009), and use DP-SGD (Abadi et al., 2016) to optimise its variational parameters and hyperparameters. This is a strong baseline because GPs excel in small data, and DP-SGD is a state-of-the-art DP fine-tuning algorithm. We found it critical to carefully tune the DP-SGD parameters and the GP initialisation using BayesOpt, and devoted substantial compute on this to ensure we have maximised GP performance. We refer to this baseline as the DP-SVGP. For details see Appendix D.
@@ 5.1 Synthetic tasks
DPConvCNP competes with DP-SVGP. Even in the Gaussian setting, where the DP-SVGP is given the covariance of the generative process, the DPConvCNP remains competitive (red and purple in Figure 5; top). While the DP-SVGP outperforms the DPConvCNP for some $N$ and $\ell,$ the gaps are typically small. In contrast, the DP-SVGP often fails to provide sensible predictions (see $\ell=0.25,N\geq 300$), and tends to overestimate the lengthscale, which is a known challenge in variational GPs (Bauer et al., 2016). We also found that the DP-SVGP also tends to underestimate the observation noise resulting in over-smoothed and over-confident predictions which lead to a counter-intuitive reduction in performance as $N$ increases. By contrast, the DPConvCNP gracefully handles different $N$ and recovers predictions that are close to the non-DP Bayesian posterior for modest $\epsilon$ and $N,$ with runtimes several orders of magnitude faster than the DP-SVGP (Figure 6).
@@ 6 Limitations & Conclusion
Limitations. The DPConvCNP does not model dependencies between target outputs, which is a major limitation. This could be achieved straightforwardly by extending our approach to LNPs, GNPs, or ARNPs. Another limitation is that the efficacy of any sim-to-real scheme is limited by the quality of the simulated data. If the real and the simulated data differ substantially, then sim-to-real transfer has little hope of working. This can be mitigated by simulating diverse datasets to ensure the real data are in the training distribution. However, as simulator diversity increases, predictions also typically become less certain, so there is a sweet spot in simulator diversity. While we observed strong sim-to-real results, exploring the effect of this diversity is a valuable direction for future work.
@@ Acknowledgements
This work was supported by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI as well as Grants 356499 and 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. SM is supported by the Vice Chancellor’s and Marie and George Vergottis Scholarship, and the Qualcomm Innovation Fellowship. RET is supported by Google, Amazon, ARM, Improbable and EPSRC grant EP/T005386/1.

### v2

@@ 1 Introduction
Neural Processes. Our method is based on neural processes (NPs; Garnelo et al., 2018a), a model which leverages the flexibility of neural networks to produce well-calibrated predictions in the meta-learning setting. The parameters of the NP are meta-trained to generalise to unseen datasets, while adapting to new contexts much faster than gradient-based fine-tuning alternatives (Finn et al., 2017).
@@
  2. 2.
    We improve upon the functional mechanism of Hall et al. (2013) by leveraging Gaussian DP theory (Dong et al., 2022), showing that context set privacy can be protected with smaller amounts of noise (at least 25% lower standard deviation in the settings considered in Figure 4). We incorporate these improvements into DPConvCNP, but note that they are also of interest in any use case of the functional mechanism.
  3. 3.
@@ 2 Related Work
Training deep learning models on public proxy datasets and then fine-tuning with DP on private data is becoming increasingly common in computer vision and natural language processing applications (Yu et al., 2021; Li et al., 2022; De et al., 2022; Tobaben et al., 2023). However, these approaches rely on the availability of very large non-sensitive datasets. Because these datasets would likely need to be scraped from the internet, it is unclear whether they are actually non-sensitive (Tramèr et al., 2024). On the other hand, other approaches study meta-learning with DP during meta-training (Li et al., 2020; Zhou and Bassily, 2022), but do not enforce privacy guarantees at meta-test time.
@@ 3.1 Meta-learning and Neural Processes
In CNPs, a standard choice, which we also use here, is to let $\pi_{\phi,\theta}(\mathbf{x}^{(t)},D^{(c)})$ return a mean $\mu_{\phi,\theta}(x^{(t)},D^{(c)})$ and a variance $\sigma_{\phi,\theta}^{2}(x^{(t)},D^{(c)}),$ to parameterise a predictive distribution that factorises across the target points $y^{(t)}|x^{(t)}\sim\mathcal{N}(\mu_{\phi,\theta}(x^{(t)},D^{(c)}),\sigma_{\phi,\theta}^{2}(x^{(t)},D^{(c)})).$ We note that our framework straightforwardly extends to more complicated $\pi_{\phi,\theta}(\mathbf{x}^{(t)},D^{(c)})$. To train a CNP to make accurate predictions, we can optimise a log-likelihood objective Garnelo et al. (2018a) such as
@@
  Input: Simulated datasets $(D_{m})^{M}_{m=1}$, encoder $\texttt{enc}_{\phi}$, decoder $\texttt{dec}_{\theta},$ iterations $T$, optimiser opt
@@
  Input: Real context $D^{(c)}$, $\texttt{enc}_{\phi}$, $\texttt{dec}_{\theta}$
@@ 3.2 Differential Privacy
We consider $D\in\mathbb{R}^{N\times d}$ with $N$ users and $d$ dimensions, and use the substitution neighbourhood relation $\sim_{S}$ where $D\sim_{S}D^{\prime}$ if $D$ and $D^{\prime}$ differ by at most one row.
@@
Properties of (G)DP. Differential privacy has several useful properties. First, post-processing immunity guarantees that post-processing the result of a DP algorithm does not cause privacy loss: {theorem}[Dwork and Roth 2014] Let $\mathcal{M}$ be an $(\epsilon,\delta)$-DP (or $\mu$-GDP) algorithm and let $f$ be any, possibly randomised, function. Then $f\circ\mathcal{M}$ is $(\epsilon,\delta)$-DP (or $\mu$-GDP). Composition of DP mechanisms refers to running multiple mechanisms on the same data. When each mechanism can depend on the outputs of the previous mechanisms, the composition is called adaptive. GDP is particularly appealing because it has a simple and tight composition formula: {theorem}[Dong et al. 2022] The adaptive composition of $T$ mechanisms that are $\mu_{i}$-GDP ($i=1,\dots,T$), is $\mu$-GDP with $\mu=\sqrt{\mu_{1}^{2}+\dotsb+\mu_{T}^{2}}.$ Gaussian mechanism. One of the central mechanisms to guarantee DP, is the Gaussian mechanism. This releases the output of a function $f$ with added Gaussian noise
$\mathcal{M}(D)=f(D)+\mathcal{N}(0,\sigma^{2}I),$ (6)
where the variance $\sigma^{2}$ depends on the $l_{2}$-sensitivity of $f,$ defined as
$\Delta=\sup_{D\sim D^{\prime}}||f(D)-f(D^{\prime})||_{2}.$ (7)
{theorem}
[Dong et al. 2022] The Gaussian mechanism with variance $\sigma^{2}=\nicefrac{{\Delta^{2}}}{{\mu^{2}}}$ is $\mu$-GDP.
@@ 3.3 The Functional Mechanism
The functional mechanism with multiplier $c$ and sensitivity $\Delta$ is defined as
$\mathcal{M}(D)=f_{D}+cg.$ (9)
@@ 4 Differential privacy for the ConvCNP
Now we turn to our main contributions. First, we tighten the functional mechanism privacy analysis in Section 4.1 and then we build the functional mechanism into the ConvCNP in Section 4.2.
@@ 4.1 Improving the Functional Mechanism
Figure 4: Noise magnitude comparison for the classical functional mechanism of Hall et al. (2013), the RDP-based mechanism of Jiang et al. (2023) and our improved GDP-based mechanism. The line for Hall et al. cuts off at $\epsilon=1$ since their bound has only been proven for $\epsilon\leq 1$. We set $\Delta^{2}=10$ and $\delta=10^{-3}$, which are representative values from our experiments. See Appendix A.6 for more details.
@@ 4.2 Differentially Private Convolutional CNP
which are the two quantities we want to release under DP. To achieve this, we must first determine the sensitivity of $r^{(d)}$ and $r^{(s)},$ as defined in Eq. 8. Recall that we use the substitution neighbourhood relation $\sim_{S}$, defined as $D^{(c)}_{1}\sim_{S}D^{(c)}_{2}$ if $D^{(c)}_{1}$ and $D^{(c)}_{2}$ differ in at most one row, i.e. by a single context point. Since the RBF $\psi$ is bounded above by $1,$ it can be shown (see Section A.4) that the squared $l_{2}$-sensitivity of $r^{(d)}$ is bounded above by $2,$ and this bound is tight. Unfortunately however, since the signal channel $r^{(s)}$ depends linearly on each $y^{(c)}_{n}$ (see Eq. 10), its sensitivity is unbounded. To address this, we clip each $y^{(c)}_{n}$ by a threshold $C,$ which is a standard way to ensure the sensitivity is bounded. With this modification we obtain the following tight sensitivities for $r^{(d)}$ and $r^{(s)}$:
$\Delta_{\mathcal{H}}^{2}r^{(d)}=2,~{}~{}~{}~{}\Delta_{\mathcal{H}}^{2}r^{(s)}=4C^{2}$ (11)
With these in place, we can state our privacy guarantee which forms the basis of the DPConvCNP. Post-processing immunity (Section 3.2) ensures that post-processing $r^{(s)}$ and $r^{(d)}$ with the ConvCNP decoder does not result in further privacy loss. {theorem} Let $g_{d}$ and $g_{s}$ be sample paths of two independent Gaussian processes having zero mean and covariance function $k,$ such that $0\leq k\leq C_{k}$ for some $C_{k}>0.$ Let $\Delta_{d}^{2}=2C_{k}$ and $\Delta_{s}^{2}=4C^{2}C_{k}.$ Then releasing $r^{(d)}+\sigma_{d}g_{d}$ and $r^{(s)}+\sigma_{s}g_{s}$ is $\mu$-GDP with $\mu=\sqrt{\nicefrac{{\Delta_{s}^{2}}}{{\sigma_{s}^{2}}}+\nicefrac{{\Delta_{d}^{2}}}{{\sigma_{d}^{2}}}}.$
@@
{corollary}
Algorithm 2 with the DPSetConv encoder from Algorithm 3 is $(\epsilon,\delta)$-DP with respect to the real context set $D^{(c)}$.
Proof.
The noise_scales method in Algorithm 3 computes the appropriate $\sigma_{d}$ and $\sigma_{s}$ values from Section 4.2 and Section 3.2 such that releasing the functional encodings $r^{(d)}+\sigma_{d}g_{d}$ and $r^{(s)}+\sigma_{s}g_{s}$ is $(\epsilon,\delta)$-DP. The $(\epsilon,\delta)$-DP guarantee extends (Hall et al., 2013, Proposition 5) to the point evaluations $\mathbf{r}^{(d)}$ and $\mathbf{r}^{(s)}$ over the grid $\mathbf{x}$ in Algorithm 3. Post-processing immunity (Section 3.2) extends $(\epsilon,\delta)$-DP to Algorithm 2. ∎
@@ 4.3 Training the DPConvCNP
Privacy hyperparameters. By Section 3.2 and Section 4.2, each $(\epsilon,\delta)$-budget implies a $\mu$-budget, placing a constraint on the sensitivities and noise magnitudes, namely $\mu^{2}=\nicefrac{{\Delta_{s}^{2}}}{{\sigma_{s}^{2}}}+\nicefrac{{\Delta_{d}^{2}}}{{\sigma_{d}^{2}}}.$ Since $\psi$ is an RBF, $\Delta_{d}^{2}=2$ and $\Delta_{s}^{2}=4C^{2},$ and we need to specify $C,\sigma_{s}$ and $\sigma_{d}$, subject to this constraint. We introduce a variable $10<t<$ 1 and rewrite the constraint as
$\sigma_{s}^{2}=\frac{4C^{2}}{t\mu^{2}}~{}~{}\text{ and }~{}~{}\sigma_{d}^{2}=\frac{2}{(1-t)\mu^{2}}$ (12)
allowing us to freely set $t$ and $C.$ One straightforward approach is to fix $t$ and $C$ to hand-picked values, but this is sub-optimal since the optimal values depend on $\mu,N,$ and the data statistics. Instead, we can make them adaptive, letting $t:\mathbb{R}^{+}\times\mathbb{N}\to(0,1)$ and $C:\mathbb{R}^{+}\times\mathbb{N}\to\mathbb{R}^{+}$ be learnable functions, e.g. neural networks $t(\mu,N)=\texttt{sig}(\text{NN}_{t}(\mu,N))$ and $C(\mu,N)=\text{exp}(\text{NN}_{C}(\mu,N))$ where sig is the sigmoid. These networks are meta-trained along with all other parameters of the DPConvCNP.
@@ 5 Experiments & Discussion
We conduct experiments on synthetic and a sim-to-real task with real data. We provide the exact experimental details in Appendix E. We make our implementation of the DPConvCNP public in the repository https://github.com/cambridge-mlg/dpconvcnp.
DP-SGD baseline. Since, we are interested in the small-data regime, i.e. a few hundred datapoints per task, we turn to Gaussian processes (GP; Rasmussen and Williams, 2006), the gold-standard model for well-calibrated predictions in this setting. To enforce DP, we make the GP variational (Titsias, 2009), and use DP-SGD (Abadi et al., 2016) to optimise its variational parameters and hyperparameters. This is a strong baseline because GPs excel in small data, and DP-SGD is a state-of-the-art DP fine-tuning algorithm. We found it critical to carefully tune the DP-SGD parameters and the GP initialisation using BayesOpt, and devoted substantial compute on this to ensure we have maximised GP performance. We refer to this baseline as the DP-SVGP. For details see Appendix D.
@@ 5.1 Synthetic tasks
DPConvCNP competes with DP-SVGP. Even in the Gaussian setting, where the DP-SVGP is given the covariance of the generative process, the DPConvCNP remains competitive (red and purple in Figure 6; top). While the DP-SVGP outperforms the DPConvCNP for some $N$ and $\ell,$ the gaps are typically small. In contrast, the DP-SVGP often fails to provide sensible predictions (see $\ell=0.25,N\geq 300$), and tends to overestimate the lengthscale, which is a known challenge in variational GPs (Bauer et al., 2016). We also found that the DP-SVGP tends to underestimate the observation noise, resulting in over-smoothed and over-confident predictions which lead to a counter-intuitive reduction in performance as $N$ increases. By contrast, the DPConvCNP gracefully handles different $N$ and recovers predictions that are close to the non-DP Bayesian posterior for modest $\epsilon$ and $N,$ with runtimes several orders of magnitude faster than the DP-SVGP (Figure 5).
@@ 6 Limitations & Conclusion
Limitations. The DPConvCNP does not model dependencies between target outputs, which is a major limitation. This could be achieved straightforwardly by extending our approach to LNPs, GNPs, or ARNPs. Another limitation is that the efficacy of any sim-to-real scheme is limited by the quality of the simulated data. If the real and the simulated data differ substantially, then sim-to-real transfer has little hope of working. This can be mitigated by simulating diverse datasets to ensure the real data are in the training distribution. However, as simulator diversity increases, predictions typically become less certain, so there is a sweet spot in simulator diversity. While we observed strong sim-to-real results, exploring the effect of this diversity is a valuable direction for future work.
@@ Acknowledgements
This work was supported in part by the Research Council of Finland (Flagship programme: Finnish Center for Artificial Intelligence, FCAI as well as Grants 356499 and 359111), the Strategic Research Council at the Research Council of Finland (Grant 358247) as well as the European Union (Project 101070617). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. SM is supported by the Vice Chancellor’s and Marie and George Vergottis Scholarship, and the Qualcomm Innovation Fellowship. Richard E. Turner is supported by Google, Amazon, ARM, Improbable, EPSRC grant EP/T005386/1, and the EPSRC Probabilistic AI Hub (ProbAI, EP/Y028783/1).